OpenStudy (anonymous):

How do I integrate this? (1/ (e^x+1)) du. It looks simple but I didn't do integrals for years and forgot how to do them. Is there a rule for this?

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ 1 }{ e^x+1 } dx\]

OpenStudy (anonymous):

oh nvm u sub and partial fraction did the trick

OpenStudy (sirm3d):

\[\large \frac{1}{e^x+1}\cdot\frac{e^{-x}}{e^{-x}}=\frac{e^{-x}}{1+e^{-x}}\] substitution simplifies the integral.